this is in java program, I need to know how can I solve this in hand with out using computer program.  

public class JavaApplicationGuess {
public static void main(String[] args) {
int i = 0;
int x = 0;
  
for (i =3; i > 1; i--) {
do {
x = x * 2 + 3;
System.out.println(x);
} while (x < 8);
x = x -1 + i;
System.out.println(i + " " + x);
}
}
}

my output doing in hand are

3

9

21 which is wrong

2 22 wrong
Answer #1

ANSWER:-

initial value of i=0 and x=0

first exicute for loop and initialize i=3 then check condition , condition is true so exicute body of for loop and exicute do- while loop and exicute statement x=x*2+3 = 0*2+3=3 so value of x will be 3 and print 3,after print 3 check while condition , condition is true so exicute do-while loop again and exicute statement x=x*2+3 = 3*2+3=9 so value of x will be 9 and print 9,after print 9 check while condition,condition is false ,so terminate do-while loop and exicute satement x = x -1 + i=9-1+3=11 so value of x will be 11 then exicute print statement and print 3 11 after first iteration of for loop i will be 2 and x will be 11

and again for loop exicute for i=2 then check condition , condition is true so exicute body of for loop and exicute do- while loop and exicute statement x=x*2+3 = 11*2+3=25 so value of x will be 25 and print 25,after print 25 check while condition ,condition is false ,so terminate do-while loop and exicute satement x = x -1 + i=25-1+2=26 so value of x will be 26 then exicute print statement and print 2 26 after second iteration of for loop i will be 1 and x will be 26

and again for loop exicute for i=1 then check condition , condition is false so loop terminate and program will be end

so final output will be

3

9

3 11

25